# Data Driven Drift Correction For Complex Optical Systems

### v1

@@ 4 Conclusions
Considering the broader impact of this investigation, transitioning from feedback loops for drift correction to Time Varying Bayesian Optimization represents a fundamental paradigm shift from reactive stabilization to proactive, continuous performance optimization. Current drift correction approaches focus on correcting deviations from a pre-determined setpoint. TVBO continuously seeks an optimum in the presence of drift. In this regard, the beamline would not just be stable, but would attempt to operate at its peak achievable performance at all times, even as thermal, mechanical, and electronic conditions keep changing. Furthermore, TVBO can be used in multi-objective settings using different objective functions and constraints, which is not simple using feedback loops. These changes represent a step towards self-driving accelerator facilities, where control transitions from a system of manually-tuned, independent feedback loops to an automated agent that can re-optimize over the complex, high-dimensional and time-varying input space.
Acknowledgment Use of the Linac Coherent Light Source (LCLS), SLAC National Accelerator Laboratory, is supported by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences under Contract No. DE-AC02-76SF00515. This work was performed and partially supported by the US Department of Energy (DOE), Office of Science, Office of Basic Energy Sciences Data, Artificial Intelligence and Machine Learning at the DOE Scientific User Facilities program under the MLExchange Project (award No. 107514).
Disclosures The authors declare no conflicts of interest.
Data availability Data underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon reasonable request.

### v2

@@ 4 Conclusions
Considering the broader impact of this investigation, transitioning from human-in-the-loop drift correction to Time Varying Bayesian Optimization represents a paradigm shift from reactive stabilization to proactive, continuous performance optimization. Current drift correction approaches focus on Beam Position Error only, whereas TVBO continuously seeks an optimum in the presence of drift. In this regard, the beamline would not just be stable, but would attempt to operate at its peak achievable performance at all times, even as thermal, mechanical, and electronic conditions keep changing. Furthermore, TVBO can be used in multi-objective settings using different objective functions and constraints, which is not as simple while using feedback loops. These changes represent a step towards self-driving accelerator facilities, where control transitions from a system of manually-tuned, independent feedback loops to an automated agent that can re-optimize over the complex, high-dimensional and time-varying input space.
Use of the Linac Coherent Light Source (LCLS), SLAC National Accelerator Laboratory, is supported by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences under Contract No. DE-AC02-76SF00515. This work was performed and partially supported by the US Department of Energy (DOE), Office of Science, Office of Basic Energy Sciences Data, Artificial Intelligence and Machine Learning at the DOE Scientific User Facilities program under the MLExchange Project (award No. 107514).
The authors declare no conflicts of interest.
Data underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon reasonable request.